question 1
make use of the following quadratic equation and answer the question that follow
y=-(x-2)(x+4)
(a) what are the roots of this quadratic function?give the co-oridinates of the roots without any calulations.
(b)what is the Y-intercept of this quadraic function?give your answer in co-ordinate form-show all your working
(c) what is the equation of the axis of symmetry of this function ?Use only the roots in (a) to calculate the axis of symmetry -show all yor working
(d) What is the co-oridinate form of the Vertex for the givin quadratic function? use only your answer in (c) to calculate the vertex -show all you working .state if the vertex is minimum or maximum point-give a reason for your answer .give the minimum or maximum value -state which it is.
(e) now draw your graph ;clearly naming and showing the following .axes.axis of symmetry ,roots ,Y-intercept and vertex

Answer by Alan3354(69443)

make use of the following quadratic equation and answer the question that follow
y=-(x-2)(x+4)
(a) what are the roots of this quadratic function?give the co-oridinates of the roots without any calulations.
x = 2 --> (2,0)
x = -4 --> (-4,0)
-----------------
(b)what is the Y-intercept of this quadraic function?give your answer in co-ordinate form-show all your working
f(x) = -x^2 - 2x + 8
f(0) = 8 = y-ntercept
--------------------------
(c) what is the equation of the axis of symmetry of this function ?Use only the roots in (a) to calculate the axis of symmetry -show all yor working
Axis of symmetry = average of the x-intercepts
x = (2 - 4)/2 = -1
x = -1
------------
(d) What is the co-oridinate form of the Vertex for the givin quadratic function? use only your answer in (c) to calculate the vertex -show all you work.
f(-1) = -1 + 2 + 8 = 9
Vertex at (-1,9)
--------------------
state if the vertex is minimum or maximum point-give a reason for your answer.
It's a max. The x^2 term is negative.
---------------
give the minimum or maximum value -state which it is.
The max is the vertex, at (-1,9)
-------------------------------
(e) now draw your graph ;clearly naming and showing the following .axes.axis of symmetry ,roots ,Y-intercept and vertex
I did.